# Transformer brain encoders explain human high-level visual responses

### v1

@@ 4.2 Ensemble
To address this concern further, we consider a robust phenomenon shown consistently using linear encoding with convolutional neural network backbones, where the earlier layers of the network are better features for predicting the earlier visual areas [42, 14, 29, 17, 43]. We trained different transformer decoders with image tokens coming from different layers of the DINOv2 backbone. We then use a softmax operation across the ensemble of models to get the final prediction for each voxel. The softmax weights are based on goodness of the prediction for each model for that vertex in the validation set. Figure. 4A shows the accuracy of the overall model on the brain surface for subject 1. The layers that had the highest weights in the ensemble for predicting for each given voxel is shown in  4B; higher visual areas were better predicted by later encoder layers, indicating that encoder layers capture similar feature abstractions as the brain.
@@ 4.3 Attention maps
Figure 5: Attention maps. Transformer decoder cross attention scores for three ROIs overlaid on the images. The selected ROIs show different ways in which the learned ROI queries can route information— based on location (V2d), content (FBA), or a combination of the two (OFA) depending on the location of the ROI in the brain processing hierarchy.
@@
The difference in our approach is that the cross-attention scores (Fig. 1B) can be examined to reveal the selectivity for each ROI making our model inherently more interpretable. We visualize the attention maps for 3 different ROIs in Figure 5 for the transformer encoder trained with ROI decoder queries with DINOv2 backbone. First is an early visual area, V2d (dorsal) in the left hemisphere. Since the visual field is flipped around both horizontal and vertical meridians in the cortex (starting from the retina), we expect the brain activity in this area to represent visual information from the bottom-right of the input (given that the subjects were instructed to hold fixation at the center of the screen for the presentation duration). We see this exact pattern emerge in the attention maps. Recall that the decoder queries can learn to attend to both patch locations or their content (since the key value is the sum of backbone image patches and positional encoding). In this case, the attention seems to completely be driven by the location, similarly for all the images, ignoring the content. This is exactly what we would expect from an early visual area. The fact that all the vertices in this ROI have to share the same attention map hurts accuracy as we saw in Figure  2D since the vertices do have smaller RFs in this area than a quadrant, however this can be addressed by vertex level routing.
The second ROI is OFA in the right hemisphere, a mid-level face selective area [12]. The attention maps is this ares consistently focus on faces. Since this area is in the right hemisphere it also has a preference for visual input in the left visual field. We can see this for cases with multiple faces where the second face in the right visual field in not strongly attended. The decoder query therefore makes use of both the positional encoding and he content component of the key to attend to the most relevant part of the image to predict vertices in this ROI. The attention could also be spread across multiple faces in different locations. This is the important dynamic aspect of the receptive field in higher visual areas that can be captured using the transformer attention mechanism. The third area is FBA in the right hemisphere, a high level body selective area [31]. The attention maps are more spread across bodies for this ROI and not just faces. In the Supplementary section A.2, we provide an analyses of the similarity between the learned queries for different ROIs (capturing visual and semantic similarity between them) and also show how our model can be used in an interpretability pipeline using diffusion models [24] to generate stimuli that maximally activate different ROIs (section A.3).
@@ 4.4 Text modality
We have tested the transformer encoding model on a few vision backbones but it remains to test whether this approach is generalizable to other modalities. TO test this, we first used the BLIP model [23] to generate short captions for all the images in the dataset. Using BERT [8] as the feature backbone, the decoder work exactly as before, using ROI queries to map backbone features to fMRI responses. Table 5 shows how the transformer model outperforms the regression model across all subjects (with a fraction of the parameters). Given only semantic information available in the captions, the model can only predict the high level visual areas as shown in Figures  6A and  6B.
@@ 5 Discussion
Linear encoding models have been the dominant method used for learning the mapping from model features to brain activity [10]. The reasons for this (see [16] for a review of these points) include theoretical simplicity, allowing comparison among backbone features, biological plausibility, and the ability to interpret the learned weights. However, this approach is parameters inefficient for a typical number of voxels and image features, ignores the organization of the features, and does not capture nonlinear computations between brain areas such as ubiquitous normalizations [38]. Our proposed routing based method not only reaches state of the art accuracy, it also achieves the aforementioned desiderata for encoding models, as we have shown in our results.
Foundation models (e.g. DONOv2 or CLIP for vision) trained with self-supervised objectives can serve as general visual representation backbones. However these task agnostic models do not capture all the computations in the brain and between brain areas, which needs be addressed by learning better encoding models. Our work suggests a mechanism for how different brain areas dynamically gate their input based on the input content and the area selectivity. Our results showing that the encoding accuracy for high-level areas cannot be improved beyond ROI-based routing also agrees with prior work on between area interactions using communication subspaces [37]. The routed information that is relevant to an area can then get expanded more in-depth. This process allows for cutting down on wiring cost in the brain by not connecting all the units in one area to another area but rather only a subset of relevant information getting routed with more local connections expanding the representation.
@@
We performed our experiments on NSD [2], the largest image viewing fMRI dataset to date. It will be important to test the generality of our approach on other datasets using different recording techniques (Neurophysiology, EEG, etc) and on different input modalities (such as video and audio). We used vertex-wise routing to capture the responses in early visual areas but while the computations for smaller receptive fields can be learned by this approach, the way the RFs are implemented in the brain are through different anatomical and wiring constraints. Also we chose for the model to read out the brain responses from a backbone for both early and high-level visual areas. Future work will seek to explore the connectivity between early and high-level visual areas in a more integrated system and test whether making the model further aligned with known anatomy of the visual cortex will improve performance.
@@ A.1 Encoding accuracies for Subjects 2, 3 and 7
Figure S2: The differences in encoding accuracy between the transformer and the ridge regression models showing that any improvement in the former is driven by better prediction of higher visual areas.
@@
A.2 Analyzing learned ROI queries
@@
A.3 Generating maximally activating images for ROIs

### v2

@@ 4.2 Ensemble
To address this concern further, we consider a robust phenomenon shown consistently using linear encoding with convolutional neural network backbones, where the earlier layers of the network are better features for predicting the earlier visual areas [52, 19, 37, 23, 53]. We trained different transformer decoders with image tokens coming from different layers of the DINOv2 backbone. We then use a softmax operation across the ensemble of models to get the final prediction for each voxel. The softmax weights are based on goodness of the prediction for each model for that vertex in the validation set. Figure. 4A shows the accuracy of the overall model on the brain surface for subject 1. The layers that had the highest weights in the ensemble for predicting for each given voxel is shown in  4B; higher visual areas were better predicted by later backbone layers, indicating that backbone layers capture similar feature abstractions as the brain.
@@ 4.3 Attention maps
Figure 5: Attention maps. Transformer decoder cross attention scores for three ROIs overlaid on the images (with brighter colors indicating higher attention weights). The selected ROIs show different ways in which the learned ROI queries can route information— based on location (V2d), content (FBA), or a combination of the two (OFA) depending on the location of the ROI in the brain processing hierarchy.
@@
The difference in our approach is that the cross-attention scores (Fig. 1B) can be examined to reveal the selectivity for each ROI, making our model inherently more interpretable. We visualize the attention maps for 3 different ROIs in Figure 5 for the transformer encoder trained with ROI decoder queries with DINOv2 backbone. First is an early visual area, V2d (dorsal) in the left hemisphere. Since the visual field is flipped around both horizontal and vertical meridians in the cortex (starting from the retina), we expect the brain activity in this area to represent visual information from the bottom-right of the input (given that the subjects were instructed to hold fixation at the center of the screen for the presentation duration). We see this exact pattern emerge in the attention maps. Recall that the decoder queries can learn to attend to both patch locations or their content (since the key value is the sum of backbone image patches and positional encoding). In this case, the attention seems to be completely driven by the location, similarly for all the images, ignoring the content. This is exactly what we would expect from an early visual area. The fact that all the vertices in this ROI have to share the same attention map hurts accuracy as we saw in Figure  2D, since the vertices do have smaller RFs in this area than a quadrant, however this can be addressed by vertex level routing.
The second ROI is OFA in the right hemisphere, a mid-level face selective area [16]. The attention maps is this area consistently focus on faces. Since this area is in the right hemisphere it also has a preference for visual input in the left visual field. We can see this for cases with multiple faces where the faces in the right visual field are not strongly attended. The decoder query therefore makes use of both the positional encoding and the content component of the key to attend to the most relevant part of the image to predict vertices in this ROI. The attention could also be spread across multiple faces in different locations. This is the important dynamic aspect of the receptive field in higher visual areas that can be captured using the transformer attention mechanism. The third area is FBA in the right hemisphere, a high level body selective area [39]. The attention maps are more spread across bodies for this ROI and not just faces. Supplementary section A.2 includes a quantitative analyses of the category selectivity of the attention maps. In the Supplementary section A.3, we provide an analyses of the similarity between the learned queries for different ROIs (capturing visual and semantic similarity between them) and also show how our model can be used in a pipeline using diffusion models [32] to generate stimuli that maximally activate different ROIs (section A.4).
@@ 4.4 Text modality
So far we have tested the transformer encoding model on a few vision backbones but is this approach generalizable to other modalities? To test this, we first used the BLIP model [31] to generate short captions for all the images in the dataset. Using BERT [10] as the feature backbone, the decoder works exactly as before, using ROI queries to map backbone features to fMRI responses. Table 6 shows how the transformer model outperforms the regression model across all subjects (with a fraction of the parameters). Given only semantic information available in the captions, the model can only predict the high level visual areas as shown in Figures  6A and  6B.
@@ 5 Discussion
Linear encoding models have been the dominant method used for learning the mapping from model features to brain activity [14]. The reasons for this (see [21] for a review of these points) include theoretical simplicity, allowing comparison among backbone features, biological plausibility, and the ability to interpret the learned weights. However, this approach is parameter inefficient for a typical number of voxels and image features, ignores the organization of the features, and does not capture nonlinear computations between brain areas such as ubiquitous normalizations [48]. Our proposed routing based method not only reaches state of the art accuracy, it also achieves the aforementioned desiderata for encoding models, as we have shown in our results.
Foundation vision models (e.g. DONOv2 or CLIP) trained with self-supervised objectives can serve as general visual representation backbones. However these task agnostic models do not capture all the computations in the brain and between brain areas, which needs to be addressed by learning better encoding models. Our work suggests a mechanism for how different brain areas dynamically gate their input based on the input content and the area selectivity. A flexible routing mechanism is reflected in deviation from the classical RF characterization of responses, so content-dependent RF shifts provide evidence for a more flexible mechanism [42, 18]. Our results showing that the encoding accuracy for high-level areas cannot be improved beyond ROI-based routing also agrees with prior work on between area interactions using communication subspaces [47] which can also be modulated by attention. The routed information that is relevant to an area can then get expanded more in-depth. This process allows for cutting down on wiring cost in the brain by not connecting all the units in one area to another area but rather only a subset of relevant information getting routed with more local connections expanding the representation.
@@
We performed our experiments on NSD [3], the largest image viewing fMRI dataset to date. It will be important to test the generality of our approach on other datasets using different recording techniques (Neurophysiology, EEG, etc) [13] and on different input modalities (such as video and audio). We used vertex-wise routing to capture the responses in early visual areas but while the computations for smaller receptive fields can be learned by this approach, the way the RFs are implemented in the brain are through different anatomical and wiring constraints. Also we chose for the model to read out the brain responses from a backbone for both early and high-level visual areas. Future work will seek to explore the connectivity between early and high-level visual areas in a more integrated system and test whether making the model further aligned with known anatomy of the visual cortex will improve performance.
Acknowledgments
Research reported in this publication was supported in part by the National Institute of Neurological Disorders and Stroke of the National Institutes of Health under award numbers 1RF1NS128897 and 4R01NS128897. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ A.1 Encoding accuracies for Subjects 2, 3 and 7
Figure S2: The differences in encoding accuracy between the transformer and the ridge regression models shows that the transformer encoder better predicts especially higher visual areas.
@@
A.2 Category selectivity of attention maps
To quantify the category selectivity of attention maps, we classified each pixel of the test set images using YOLOv5 [22] and YOLOv8-face [11] into five categories: background, face, body (pixels classified as person but not as face), animal, and food. For each ROI, we calculated and resized its attention maps to $434\times 434$ for these images, and reported the categories of the 2k pixels with top attention values. We found that the category selectivity is consistent with ROI labels, with EBA most selective for body, FFA most selective for face, and OPA/PPA/RSC most selective for background.
Table 7: Category selectivity of ROI attention for subject 1.
background face body animal food
EBA 0.03 0.36 0.61 0.00 0.00
FFA-1 0.00 0.79 0.16 0.05 0.00
FFA-2 0.00 0.83 0.17 0.00 0.00
OPA 0.54 0.12 0.14 0.05 0.15
PPA 0.44 0.25 0.11 0.10 0.10
RSC 0.66 0.23 0.11 0.00 0.00
Table 8: Category selectivity of ROI attention for subject 2.
background face body animal food
EBA 0.03 0.25 0.72 0.00 0.00
FFA-1 0.05 0.57 0.19 0.18 0.00
FFA-2 0.05 0.53 0.27 0.15 0.00
OPA 0.74 0.15 0.06 0.05 0.00
PPA 0.78 0.11 0.08 0.03 0.00
RSC 0.71 0.19 0.09 0.00 0.00
Table 9: Category selectivity of ROI attention for subject 5.
background face body animal food
EBA 0.29 0.26 0.40 0.05 0.00
FFA-1 0.12 0.67 0.13 0.08 0.00
FFA-2 0.10 0.59 0.28 0.03 0.00
OPA 0.31 0.33 0.20 0.16 0.00
PPA 0.36 0.29 0.20 0.15 0.00
RSC 0.08 0.37 0.50 0.00 0.06
Table 10: Category selectivity of ROI attention for subject 7.
background face body animal food
EBA 0.38 0.05 0.43 0.09 0.05
FFA-1 0.00 0.88 0.02 0.10 0.00
FFA-2 0.17 0.20 0.38 0.20 0.05
OPA 0.40 0.18 0.14 0.28 0.00
PPA 0.51 0.29 0.16 0.05 0.00
RSC 0.56 0.25 0.17 0.02 0.00
A.3 Analyzing learned ROI queries
@@
A.4 Generating maximally activating images for ROIs
@@
Appendix B Compute used
We used GPUs (NVIDIA L40s), memory, and storage resources from an internal cluster. Storage for the entire project totals roughly 3TB. Training the model used roughly 4,000 GPU hours. Running the remaining experiments used roughly 1,000 GPU hours. The full project required more compute than these estimates due to failed experiments, experiments not included in the paper, and model iteration.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All our claims are directly backed by presented results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations on page 9.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We include all the experimental details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will upload the code as part of the supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We do provide the training/testing details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: We provide the results for all the subjects in our dataset separately. The reason as described in the text, is that different subjects have slightly different number of vertices and ROI boundaries, so we felt it was more appropriate to not average over subjects and to just transparently report the data for all of them.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We are including all the details about computing resources in the supplementary files.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: answerYes
  44. Justification: Our work conforms with the code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the societal impacts of our work in the supplementary section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We used publicly available datasets. The encoding models do not have a high risk of misuse to our knowledge.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We used publicly available fMRI datasets that we credit. We own the other assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All the assets are well documented and the documentation is provided.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We use publicly available datasets and the original papers can be consulted for all the details. We provide the relevant information in our paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We used public available datasets.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We did not use LLMs in any significant way.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.